OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with an x-intercept of -4 and a y-intercept of -2

OpenStudy (jmark):

Given x-intercept of (-4,0) and a y-intercept of(0, -2) find the slope ,m=[y2-y1]/[x2-x1] =>m=[-2-0]/[0+4] =>m=-2/4 =>m=-1/2 then equation of slope =>y=mx+b =>y= -1/2 x-2 CBSE Maths Sample Paper for Class 10 http://goo.gl/GpOISh